1. Find grad f from the following functions 0) f =(x - 1) (4-1) (ii)f=2x2 + 5y2 (iii) f = x + ys v)f=(x2+ y2) (x2-y)
Answer #1

0 fr ) (4-4- grad (f) = 2 fa, fy> grad (f) =< 74-1, 41-47 f = 2x² +592 for 40 fy = log grad (f) = < fri fy > <40, loy? fa ust